OpenStudy (jamaljunior):

will fan and medal for help Solve for a. 15(25−5a)=4−a

TheSmartOne (thesmartone):

Hey Jamal! We need to use the distributive property first! Have you seen this property before? x(y - z) = xy - xz

OpenStudy (jamaljunior):

no

TheSmartOne (thesmartone):

So we need to use that propery and distribute the 15 in to 25 - 5a The x, y, and z are all variables and we can replace them with what we were given in the equation

OpenStudy (jamaljunior):

okay

OpenStudy (jamaljunior):

my apologies the 15 is actually 1 over 5 fraction

TheSmartOne (thesmartone):

It doesn't change anything (: 1/5(25 - 5a) = (1/5 * 25) - (1/5 * 5a) Can you simplify it more?

OpenStudy (jamaljunior):

im not sure

OpenStudy (jamaljunior):

hey